OpenStudy (anonymous):

Need HELP! standard form of x^2-4y^2-6x-7=0

OpenStudy (anonymous):

do you have any idea of what is standard form in the first place?

OpenStudy (anonymous):

i know the problem is a hyperbola, and the standard form for it is x^2 / a^2 - y^2 / b^2

OpenStudy (anonymous):

yes ur right.. it is a hyperbolic paraboloid.

OpenStudy (anonymous):

i know how it's supposed to be solved but it doesn't look right. . .

OpenStudy (anonymous):

z = y^2/b^2-x^2/a^2 \[Given: x^2-4y^2-6x-7=0\] The standard form for a hyperbolic paraboloid is [x/c=z^2/a^2-y^2/b^2]

OpenStudy (anonymous):

so is it vertical?